# OpenFactCheck: A Unified Framework for Factuality Evaluation of LLMs

### v1

@@ 3.3 CheckerEval
Automatic fact-checking systems aim to identify whether a claim or a document is factually correct or not with/without references, but the results are not necessarily correct. To assess the accuracy of automatic fact-checkers, we gather four LLM factuality benchmarks with human-annotated factual labels for three levels of granularity text: claims/segments/documents given (question, ChatGPT response) pairs, including FacTool, FELM-WK, Factcheck-GPT and HaluEval, a total of 4,835 examples, and we refer to them as FactBench. We use precision, recall, and F1-score with respect to the True and False claims and documents to evaluate the effectiveness of fact-checking systems.
This method regards the system as a whole, only assessing the final verification results, i.e., whether a claim or a document is true or false. Without the evaluation of intermediate results, it is difficult to localize which step eventually results in the erroneous factual judgment for claims Wang et al. (2023). We will incorporate the evaluation of intermediate results throughout the fact-checking pipelines in future updates.
3.4 Web Client
To enhance the user interaction experience, the corresponding UI interfaces have been developed for all functionalities within the demo. Specifically, we develop a web client based on Streamlit, which consists of four interfaces, each corresponding to one of the three modules, along with a leaderboard.
CustChecker interface in Figure 4 primarily communicates with the CustChecker backend (§3.1). Users can freely select different combinations of claim processors, retrievers, and verifiers. When given an input document or claim, the CustChecker back-end executes the fact-checking pipeline. The final verification results and the intermediate processing outcomes are then presented on the page to users for reference.
LLMEval page in Figure 5 corresponds to the module of LLMEval3.2). Users first download our predefined question set and then use their LLM for inference. After uploading the model responses, the system forwards them to background tasks, using our LLMEval for evaluation. Afterwards, a comprehensive report is generated and emailed to the user, notifying them of the availability of the report’s PDF for download. Moreover, if users consent to publish the evaluation results, we display them on the corresponding leaderboard page.
CheckerEval page in Figure 6 corresponds to the module of CheckerEval3.3), evaluating the performance of fact-checking systems. Users can download claims or documents to be checked from this page, and then use their fact-checking system to predict factuality. The results including True/False, time, and USD costs are subsequently uploaded. We evaluate the submitted fact-checker results based on the ground truth labels of the human-annotated datasets, we rank and we display them on the leaderboard.
The design principle of our web client is to invoke these functional modules in the form of third-party independent applications, without excessively intervening in the system’s architecture. Consequently, our system is made available to users in the form of a library, a command-line toolkit, and a web application.
@@ 4 Experiments
In this section, we first evaluate the performance of different models, and then we assess the accuracy of different automatic fact-checking systems in different settings.
@@
Six datasets 132.1 121.9 82.2
Table 3: Word length for responses of three LLMs over the datasets of FacTool-QA, FELM-WK, Factcheck-Bench and over six evaluation datasets except for FactScore-Bio.
@@ 4.1 LLaMA-2 and GPT-4 Evaluation
Based on questions/instructions in FactQA, we collected responses from LLaMA-2 (7B, 13B) and GPT-4. As shown in Table 3, the responses of GPT-4 tend to be shorter than that of LLaMA-2.
@@
On the Snowball dataset, we observe high error rates: >80% for LLaMA-2 and 65.5% for GPT-4, similar to the evaluation results on GPT-3.5-Turbo presented in Zhang et al. (2023a). However, when justifying previously generated content, GPT-4 can identify 87% of its own mistakes. Therefore, in these cases, errors are mostly attributed to the over-committing to the previously generated false context, rather than to large knowledge gaps in LLMs. This phenomenon is referred to as hallucination snowballing: an LLM over-commits to early mistakes, leading to more mistakes that it otherwise would not have made. Its prevalence in generative models leads to factual errors for simple facts.
Higher precision than recall is achieved on SelfAware across three models when setting a positive label as unanswerable. This reveals that many truly unanswerable questions are incorrectly recognized as answerable, implying that models are always not aware of what they do not know. Poor performance on questions with rapidly changing answers (FreshQA) illustrates the inherent challenge of retrieving up-to-date information for LLMs.
@@
  • GPT-4 has the best factuality performance with a smaller number of false claims and higher percentage of true claims, followed by LLaMA-2 13 and then 7B;
@@
FactScore LLaMA-Inst Wiki/BM25 0.87 0.74 0.80 0.34 0.56 0.42 0.82 0.68 0.74 0.34 0.52 0.41 0.76 0.66 0.71 0.34 0.46 0.39
FacTool LLaMA-Inst Web/Serper 0.88 0.80 0.84 0.40 0.56 0.47 0.93 0.38 0.54 0.32 0.91 0.47 0.79 0.31 0.44 0.30 0.79 0.44
@@
Table 5: Verification results for human-annotated claims in Factcheck-Bench. FacTool-QA, and FELM-WK. Shown are results for judging whether or not a claim is factually true or false with external knowledge (Wikipedia or Web articles) as evidence. The implementation of Factcheck-GPT uses langchain AutoGPT. LLaMA-Inst refers to LLaMA3-Instruct-8B. GPT-4 used gpt-4-turbo-2024-04-09.
@@ 4.2 Evaluating Fact-Checking Systems
We investigate automatic fact-checking systems in three aspects: accuracy, latency, and costs. Based on annotated factual labels for claims from three benchmarks including Factcheck-Bench, FacTool-QA, and FELM-WK, we evaluate the verification performance in multiple settings across different fact-checking frameworks, evidence sources, and verifiers.
Pipeline and core component modules of different fact-checking frameworks are basically similar, including obtaining atomic claims, collecting evidence and verifying correctness; all thus, while the implementations are different. For example, in terms of how to extract atomic claims, RARR does not include this step. FactScore first breaks down a document into paragraphs, and then applies NLTK (Bird and Loper, 2004) to split paragraphs into sentences, and then prompts LLMs to decompose to into atomic claims (e.g., GPT3 was used in the original implementation). However, this implementation neglects the decontextualization in the paragraph and in the sentence decomposition, making claims non-independent (e.g., He is a university professor and the CEO of a tech startup company). To mitigate, FacTool directly extracts claims based on the document, and Factcheck-GPT decontextualizes both sentences and claims based on the document.
Experimental setup:
To ensure that all checking systems verify the same sets of annotated claims with factual labels, we skip the step of extracting atomic claims from the documents. All fact-checking systems get a claim as an input, and they are expected to output whether or not the claim is true.
@@
In Table 5, we observe that automatic fact-checking systems struggle to detect false claims. Across the three datasets we experiment with, it is consistently more arduous for these systems to differentiate false claims compared to identifying true ones. This challenge may arise from the tendency of returning invalid evidence for false claims.
Retrieving evidence from the web using Serper (Google search engine results) is more effective than sourcing related passages from Wikipedia articles using BM25, given that a wider array of effective evidence is accessible on open web pages for open-domain questions. The verification accuracy of an LLM-based verifier primarily relies on the capabilities of the LLM and the effectiveness of the prompts used. For instance, the overall performance of GPT-4 surpasses that of both LLaMA-3-8B and GPT-3.5-Turbo, and thus the verification results (Factcheck-GPT) outperform those of FacTool, FactScore and Perplexity.ai, despite all of them utilizing evidence sourced from the web. While Factcheck-GPT exhibits superior effectiveness, it is associated with considerable latency and substantial costs (see Table 6).
@@ 5 Conclusion and Future Work
We proposed OpenFactCheck, a unified, easy-to-use and extensible toolkit for LLM factuality evaluation. OpenFactCheck establishes a unified framework with clearly defined modules and flexible configurations to support the customization of automatic fact-checking systems. From the application level, OpenFactCheck allows general users to check whether a claim and a document are factual or not. This proposed framework could also facilitate LLM practitioners and developers to effectively and efficiently evaluate the factuality of their LLMs from various perspectives and to assess the accuracy of automatic fact-checkers.
Our extensive experiments indicate that more than 90% of the claims generated by LLMs in response to open-domain questions are factually correct. Nevertheless, models encounter challenges when addressing some straightforward questions such as Is 7411 a prime number? This difficulty is attributed to recent findings suggesting that LLMs demonstrate weaker comprehension abilities relative to their generation capabilities. Additionally, prevalent fact-checking systems struggle to identify false claims, with the retrieval of pertinent evidence posing a significant bottleneck. The latency and the cost associated with these systems primarily hinge on implementation strategies. In the future, we will continue to integrate new techniques, features, and evaluation benchmarks to OpenFactCheck to facilitate the research progress of LLM fact-checking.
@@ A.1 FactQA Component Datasets
Snowball dataset Zhang et al. (2023a) comprises three question–answering subsets: primality testing, senator search, and graph connectivity, each with 500 yes/no questions. They aim to investigate snowballing hallucination when a model immediately outputs an incorrect answer (yes or no) as false generated context. Specifically, they prompt the language model to first output a yes/no answer and then to provide explanations. When the immediate answer is wrong, the model tends to continue to snowball the false statements instead of correcting them.
@@
Factcheck-Bench (Wang et al., 2023) Factcheck-GPT gathered a total of 94 highly challenging questions from sources including Twitter posts, internal brainstorming, and Dolly-15k, encompassing 678 claims.
@@
Domain Size
History 771
Biography 683
Mathematics 612
Transportation 519
Biology 259
Philosophy 229
Technology 208
Entertainment 191
Psychology 169
Sports 157
Science 143
Physics 136
Social Sciences 111
Literature 100
Geography 87
Astronomy 82
Economics 69
Music 66
Religion 63
General Knowledge 53
@@ A.2 Other Datasets
HaluEval Li et al. (2023a) comprises 5,000 general user queries accompanied by ChatGPT responses and 30,000 specialized examples from three distinct tasks: question answering, knowledge-grounded dialogue, and text summarization.
Self-Contradictory Anonymous (2023) constructed a dataset by sampling language model responses about various topics, consisting of 360 text descriptions covering 30 diverse topics.
@@
Appendix B Web Pages
Screenshots of four web pages are demonstrated below.
Appendix C Claim Statistics of LLM Responses
@@
Appendix D OpenAI GPT Models

### v2

@@ 3.3 CheckerEval
Automatic fact-checking systems aim to identify whether a claim or a document is factually correct or not with/without references, but the results are not necessarily correct. To assess the accuracy of automatic fact-checkers, we gather four LLM factuality benchmarks with human-annotated factual labels for three levels of granularity text: claims/segments/documents given (question, LLM response) pairs, including FacTool-QA, FELM-WK, Factcheck-Bench and HaluEval as shown in Table 2, see examples in Figures 3 and 4. We refer to them as FactBench. We highlight that FactQA is a question set without answers and human-annotated claims, while FactBench is a benchmark with annotations. We use precision, recall, and F1-score with respect to the True or False claim/document to evaluate the effectiveness of fact-checking systems.
This method regards the system as a whole, only assessing the final verification results, i.e., whether a claim or a document is true or false. The evaluation of intermediate results throughout the fact-checking pipelines will be incorporated in future updates, to localize which step eventually results in the erroneous factual judgment for claims.
@@ 4 Experiments
We first evaluate the factuality of three LLMs, and then we assess the accuracy of different automatic fact-checking systems in multiple settings.
@@
FactQA 132.1 121.9 82.2
Table 3: Word-level length for responses of LLaMA-2 7B, 13B and GPT-4 over the datasets: FacTool-QA, FELM-WK, Factcheck-Bench and FactQA.
@@ 4.1 LLaMA-2 and GPT-4 Evaluation
Based on questions/instructions in FactQA, we collected responses from LLaMA-2 (7B, 13B) and GPT-4. Default decoding parameter of openai.chat.completions.create() was used for GPT-4. Decoding parameters of LLaMA-2 7B and 13B were taken from the default generation configurations provided in the huggingface transformers library, where we set the max_token as 512 to avoid the out of memory issue. As shown in Table 3, the responses of GPT-4 tend to be shorter than that of LLaMA-2.
@@
On the Snowball dataset, we observe high error rates: >80% for LLaMA-2 and 65.5% for GPT-4, similar to the results on GPT-3.5-Turbo presented by Zhang et al. (2024). However, when justifying previously generated content, GPT-4 can identify 87% of its own mistakes. Therefore, in these cases, errors are mostly attributed to the over-committing to the previously generated false context, rather than to large knowledge gaps in LLMs. An LLM over-commits to early mistakes, leading to more mistakes that it otherwise would not have made. Its prevalence in generative models leads to factual errors for simple facts.
SelfAware aims to evaluate LLMs’ ability to understand their own limitations and unknowns, identifying unanswerable or unknowable questions. Higher precision than recall is achieved across three models with regard to unanswerable questions. This reveals that many truly unanswerable questions are incorrectly recognized as answerable, implying that models are always not aware of what they do not know. Poor performance on questions with rapidly changing answers (FreshQA) illustrates the inherent challenge of retrieving up-to-date information for LLMs.
@@
  • GPT-4 has the best factuality performance with a smaller number of false claims and higher percentage of true claims, followed by LLaMA-2 13B and then 7B;
@@
FactScore LLaMA 3-Inst 8B Wiki/BM25 0.87 0.74 0.80 0.34 0.56 0.42 0.82 0.68 0.74 0.34 0.52 0.41 0.76 0.66 0.71 0.34 0.46 0.39
FacTool LLaMA 3-Inst 8B Web/Serper 0.88 0.80 0.84 0.40 0.56 0.47 0.93 0.38 0.54 0.32 0.91 0.47 0.79 0.31 0.44 0.30 0.79 0.44
@@
Table 5: Verification results for human-annotated claims in Factcheck-Bench, FacTool-QA, and FELM-WK, judging whether or not a claim is factually true or false with external knowledge (Wikipedia or Web articles) as evidence. The implementation of Factcheck-GPT usedlangchain AutoGPT. GPT-4 refers to gpt-4-turbo-2024-04-09.
@@ 4.2 Evaluating Fact-Checking Systems
We investigate automatic fact-checking systems in three aspects: accuracy, latency, and costs. Based on annotated factual labels for claims from three benchmarks of Factcheck-Bench, FacTool-QA, and FELM-WK, we evaluate the verification performance in multiple settings across different fact-checking frameworks, evidence sources, and verifiers.
Pipeline and core component modules of different fact-checking frameworks are basically similar, including obtaining atomic claims, collecting evidence and verifying correctness; all thus, while the implementations are different. For example, in terms of how to extract atomic claims, RARR does not include this step. FactScore first breaks down a document into paragraphs, and then applies NLTK (Bird and Loper, 2004) to split paragraphs into sentences, and then prompts LLMs to decompose to atomic claims (GPT3 was used in the original paper). However, this implementation neglects the decontextualization in the paragraph and in the sentence decomposition, making claims non-independent (e.g., He is a university professor and the CEO of a tech startup company). To mitigate, FacTool directly extracts claims based on the document, and Factcheck-GPT decontextualizes both sentences and claims based on the document.
Experimental Setup
To ensure that all fact-checking systems verify the same sets of annotated claims, we skip the step of extracting atomic claims from the documents. All systems get a claim as an input, and they are expected to output whether or not the claim is true.
@@
In Table 5, we observe that automatic fact-checking systems struggle to detect false claims. Across the three datasets we experiment with, it is consistently more arduous for systems to differentiate false claims compared to identifying true ones. This challenge may arise from the tendency of returning invalid evidence for false claims.
Retrieving evidence from the web using Serper (Google search engine results) is more effective than sourcing related passages from Wikipedia articles using BM25, given that a wider array of effective evidence is accessible on open web pages for open-domain questions. The verification accuracy of an LLM-based verifier primarily relies on the capabilities of the LLM and the effectiveness of the prompts used. For instance, the overall performance of GPT-4 surpasses that of both LLaMA-3-8B and GPT-3.5-Turbo, and thus the verification results of Factcheck-GPT outperform those of FacTool, FactScore and Perplexity.ai, despite all of them utilizing evidence sourced from the web. While Factcheck-GPT exhibits superior effectiveness, it is associated with considerable latency and substantial costs (see Table 6).
@@ 5 Conclusion and Future Work
We proposed OpenFactCheck, a unified, easy-to-use and extensible framework. It supports the customization and evaluation of automatic fact-checking systems and LLM factuality evaluation. Specifically, OpenFactCheck allows general users to check whether a claim and a document are factual or not, and also facilitate LLM practitioners and developers to effectively and efficiently evaluate the factuality of their LLMs from various perspectives, and to assess the accuracy of automatic fact-checking systems.
Our extensive experiments indicate that more than 90% of the claims generated by LLMs in response to open-domain questions are factually correct. Nevertheless, models encounter challenges when addressing some straightforward questions such as Is 7411 a prime number? This difficulty can be attributed to the fact that LLMs demonstrate weaker comprehension abilities relative to their generation capabilities. Additionally, prevalent fact-checking systems struggle to identify false claims, with the retrieval of pertinent evidence posing a significant bottleneck. The latency and the cost associated with these systems primarily hinge on implementation strategies. In the future, we will continue to integrate new techniques, features, and evaluation benchmarks to OpenFactCheck to facilitate the research progress of LLM fact-checking.
@@ Acknowledgements
Limitations
While OpenFactCheck presents a comprehensive framework for factuality evaluation of LLMs, several limitations must be acknowledged:
Evaluation Datasets
The effectiveness of OpenFactCheck is dependent on the quality and diversity of the datasets used for evaluation. While we have integrated multiple datasets to cover a broad spectrum of domains and potential factual errors, the evaluation is still limited by the inherent biases and coverage gaps in these datasets. For instance, some specialized domains may not be adequately represented, potentially affecting the robustness of the evaluation for LLMs in those areas.
Moreover, we only evaluated factuality of GPT-4 and LLaMA-2 (7B, 13B) responses to demonstrate the framework’s capabilities, while the OpenFactCheck system is designed to be extensible. Future work will incorporate the new SOTA LLMs and diverse datasets to broaden the evaluation. The current evaluation provides a proof-of-concept, and we welcome collaborations to expand this further.
Latency and Costs
The performance of automatic fact-checking systems integrated within OpenFactCheck can vary significantly in terms of latency and operational costs. High accuracy often comes at the expense of increased computational resources and processing time, which may not be feasible for all users, particularly those with limited budgets or time constraints.
Reliance on External Knowledge Sources
The fact-checking modules depend heavily on external knowledge sources, such as Wikipedia and web search engines. The availability and reliability of these sources can affect the accuracy and completeness of the fact-checking process. Furthermore, the dynamic nature of web content means that the information retrieved may not always be up-to-date.
Ethical Statement
The development and deployment of OpenFactCheck are guided by a commitment to ethical principles, ensuring that the framework is used responsibly and for the benefit of society:
Transparency and Accountability
We strive to maintain transparency in the design, implementation, and evaluation of OpenFactCheck. The source code and datasets are publicly available, enabling scrutiny and fostering trust within the research community. We encourage users to report any issues or biases they encounter, facilitating continuous improvement.
Bias Mitigation
Recognizing that biases can exist in both datasets and LLMs, we are dedicated to minimizing such biases in OpenFactCheck. By integrating diverse evaluation benchmarks and encouraging the development of fair fact-checking approaches, we aim to reduce the impact of biases on factuality evaluation outcomes.
Social Impact
By enhancing the factual accuracy of LLMs, OpenFactCheck aims to contribute positively to society. Accurate information is crucial for informed decision-making and public discourse. We believe that improving the reliability of LLM outputs can help combat misinformation and support the dissemination of truthful information.
@@ A.1 FactQA Component Datasets
Snowball dataset Zhang et al. (2024) comprises three question–answering subsets: primality testing, senator search, and graph connectivity, each with 500 yes/no questions. They aim to investigate snowballing hallucination when a model immediately outputs an incorrect answer (yes or no) as false generated context. Specifically, they prompt the language model to first output a yes/no answer and then to provide explanations. When the immediate answer is wrong, the model tends to continue to snowball the false statements instead of correcting them.
@@
Factcheck-Bench (Wang et al., 2024a) Factcheck-GPT gathered a total of 94 highly challenging questions from sources including Twitter posts, internal brainstorming, and Dolly-15k, encompassing 678 claims.
@@
Domain Size Domain Size
History 771 Science 143
Biography 683 Physics 136
Mathematics 612 Social Sciences 111
Transportation 519 Literature 100
Biology 259 Geography 87
Philosophy 229 Astronomy 82
Technology 208 Economics 69
Entertainment 191 Music 66
Psychology 169 Religion 63
Sports 157 General Knowledge 53
@@ A.2 Other Datasets
HaluEval Li et al. (2023) comprises 5,000 general user queries accompanied by ChatGPT responses and 30,000 specialized examples from three distinct tasks: question answering, knowledge-grounded dialogue, and text summarization.
Self-Contradictory Mündler et al. (2024) constructed a dataset by sampling language model responses about various topics, consisting of 360 text descriptions covering 30 diverse topics.
@@
A.3 FactBench Examples
Given an example in FacTool-QA, FELM-WK and Factcheck-Bench, it includes a prompt, a response generated by GPT-3.5-turbo and human annotations: a list atomic claims decomposed based on the response, a list of factual labels [True, False, $\cdots$] corresponding to each atomic claim, and the factual label for the response, as shown in Figure 3 below:
Figure 3: An annotated FacTool-QA example sampled from FactBench.
Given an example in HaluEval, it includes a prompt, a response by GPT-3.5-turbo and human annotations: a list of false segments/spans extracted from the response, and the factual label for the response as below:
Figure 4: An annotated HaluEval example sampled from FactBench.
Appendix B Experiments
Claim Statistics of Three LLMs’ Responses
@@ Appendix B Experiments
LLM-based Verification Prompt
Figure 6 shows a prompt used to verify whether a claim is true or false based on the collected evidence.
Figure 6: LLM-based verifier prompt example.
OpenAI GPT Models
@@
Appendix C OpenFactCheck Demo
Based on this framework, we initiate a OpenFactCheck demonstration. It was implemented with a Python server, a web user interface, and a database, deployed via AWS. The Python backend can also be used as a Python toolkit, allowing easy and flexible development. This demo was further optimized and presented in Iqbal et al. (2024).
C.1 CustChecker Pseudo Code
{minted}
[…]
# FacTool claims = call_LLM(document, promot="extract context-independent atomic claims based on the document")
return claims
def retriever(claim: str, database: DB, retrieval_strategy: obj, search_api_key: str) -> List[str]: # offline DB dump evidence = retrieval_strategy(claim, database)
# online web pages by calling API evidence = serper_or_serpapi(claim, search_api_key)
return evidence
def verifier(claim: str, evidence: List[str]) -> bool: # call LLMs factual_label = call_LLM(claim, evidence, prompt="based on the evidence and your own knowledge, determine whether the claim is true or false.")
# use NLI models stance2factual = "entailment": true, "contradiction": false, "neutral": "not enough evidence" stances = [nli(evid, claim) for evid in evidence] majority_stance = majority_vote(factual_labels) factual_label = stance2factual[majority_stance]
return factual_label
Figure 7: Pseudo code for the three modules in CustChecker.
C.2 Web Client
We develop a web client based on Streamlit, consisting of four interfaces, with each corresponding to one of the three modules, along with a leaderboard, to enhance the user interaction. The design principle is to invoke these functional modules in the form of third-party applications, avoiding excessive intervention in the system’s architecture. This makes OpenFactCheck a three-in-one to users as a library, a command-line toolkit, and a web application.
As CustChecker interface shown in Figure 8, users can freely select different combinations of claim processors, retrievers, and verifiers. When given an input document or claim, the CustChecker backend (§3.1) executes the fact-checking pipeline. The final verification results and the intermediate processing outcomes are presented on the page for reference.
Figure 9 corresponds to the module of LLMEval in §3.2. Users first download our predefined question set and then upload their model responses, the system forwards them to background tasks, using LLMEval for evaluation. Afterwards, a comprehensive report is generated and emailed to the user, notifying them of the availability of the report’s PDF for download. Moreover, if users consent to publish the evaluation results, we display them on the corresponding leaderboard page.
CheckerEval page in Figure 10 evaluates the performance of fact-checking systems. Users can download claims or documents to be checked from this page, and then use their fact-checking system to predict factuality. The results including True/False, time, and USD costs are subsequently uploaded. We evaluate the submitted fact-checker results based on the ground truth labels of the human-annotated datasets, we rank and display them on the leaderboard.